# Sparse Autoencoders Learn Monosemantic Features in Vision-Language Models

### v2

@@ Abstract
Given that interpretability and steerability are crucial to AI safety, Sparse Autoencoders (SAEs) have emerged as a tool to enhance them in Large Language Models (LLMs). In this work, we extend the application of SAEs to Vision-Language Models (VLMs), such as CLIP, and introduce a comprehensive framework for evaluating monosemanticity at the neuron-level in vision representations. To ensure that our evaluation aligns with human perception, we propose a benchmark derived from a large-scale user study. Our experimental results reveal that SAEs trained on VLMs significantly enhance the monosemanticity of individual neurons, with sparsity and wide latents being the most influential factors. Notably, we demonstrate that applying SAE interventions on CLIP’s vision encoder directly steers multimodal LLM outputs (e.g., LLaVA), without any modifications to the underlying model. These findings emphasize the practicality and efficacy of SAEs as an unsupervised tool for enhancing both interpretability and control of VLMs. Code is available at https://github.com/ExplainableML/sae-for-vlm.
@@ 1 Introduction
In recent years, Vision-Language Models (VLMs) like CLIP [33] and SigLIP [45] have gained widespread adoption, owing to their capacity for simultaneous reasoning over visual and textual modalities. They have found a surge of applications in various modalities, such as in audio [10, 43] and medicine [46], transferring to new tasks with minimal supervision. However, the inner workings of these models are still not well understood. Sparse AutoEncoders (SAEs) [25] are an effective approach to probing the internal representations of such models. They efficiently discover concepts (abstract features shared between data points) through their simple architecture learned as a post-hoc reconstruction task. Although analysis with SAEs is popular for Large Language Models (LLMs) [3, 34, 17], for VLMs it has been limited to interpretable classification [35, 23], or the discovery of concepts shared across models [39].
@@
Intuitively, SAEs reconstruct activations in a higher-dimensional space to disentangle distinct concepts from their overlapping representations in neural activations [3]. Neurons at different layers within deep neural networks are known to be naturally polysemantic [29], meaning that they can fire for multiple unrelated concepts such as cellphones and rulers. One common explanation for this behavior is the superposition hypothesis [1, 13], stating that concepts are encoded as linear combination of neurons. SAEs explicitly attempt to solve this issue by separating the entangled concepts into distinct representations. Despite their widespread use in research, the absence of a metric to evaluate SAEs at the neuron-level still hinders their practicality as an interpretation tool. Discovering neurons encoding human-interpretable concepts requires analyzing them individually, which is even more tedious as both the model and SAE are wide.
@@
Figure 1: Training Sparse Autoencoders (SAEs) over VLMs (e.g. CLIP): The highly activating images of a neuron in a given layer of a pretrained VLM are polysemantic (left), and a neuron in an SAE trained to reconstruct the same layer is more monosemantic (right), reflected by higher MonoSemanticity score (MS).
@@
In this work, we quantitatively evaluate SAEs for VLMs through monosemanticity, defined as the similarity between inputs firing a neuron. We propose the MonoSemanticity score (MS) for vision tasks, that measures the pairwise similarity of images weighted by the activations for a given neuron. Unlike natural language where individual words require surrounding context (such as complete sentences) to clearly identify their meanings, individual images can directly activate neurons without additional context. We define highly activating images as images that strongly fire a particular neuron. Greater similarity among these images suggests that the neuron is more narrowly focused on a single concept, reflecting higher neuron monosemanticity. Using our MS score, we observe and validate that neurons within SAEs are significantly more monosemantic (see Figure 1, right) than the original neurons (see Figure  1, left). The neuron of the original VLM typically has a low MS score, since it fires for a wide range of objects, from cellphones to rulers. On the other hand, neurons within the SAE are more focused on a single concept, e.g. parrots, obtaining a higher score. This observation remains valid even for SAE with the same width as the original layer, implying that the sparse reconstruction objective inherently improves the separability of concepts. We further conducted a large-scale study to quantitatively assess alignment of our proposed MS score with human interpretation of monosemanticity. The results confirm that the difference between scores of two neurons is strongly correlated with humans perceiving the higher-scoring neuron as more focused on a single concept.
@@
Figure 2: Illustration of the computation of our MonoSemanticity score (MS). (A) We extract embeddings and activations from a given set of images, (B) then compute the pairwise embedding similarities and pairwise neuron activations. (C) MS is equal to the average of embedding similarities weighted by the neuron activations.
@@ 2 Related Work
Sparse Autoencoders. Recent works have focused on repurposing traditional dictionary learning methods for providing interpretability of LLMs and VLMs [2, 32]. Specifically, there has been a lot of success in interpreting and steering LLMs using the features learned by sparse autoencoders [38, 12]. With respect to the underlying SAE mechanisms, there have been several improvements proposed from the original setup, including novel activation functions such as Batch TopK [4] or JumpReLU [34], along with taking inspiration from Matryoshka representation learning [22] for SAEs [27, 5]. In this work, we provide a comprehensive analysis and evaluation of the monosemanticity at the neuron-level of SAEs in VLMs and their downstream applications.
Vision-Language Models. Since the introduction of Contrastive Language-Image Pretraining (CLIP) [33, 26, 45], several models have been introduced that either map images and text to a shared embedding space [42, 20], or develop models that are able to generate text conditioned on image inputs [24, 8]. While these models have achieved strong results on several benchmarks and use-cases over the past few years, interpretability has become an important concern, from a scientific standpoint, but also for their safe usage and scalable control. Therefore, there have been investigations into the representations of these models [28, 15, 16] which have found interpretable neurons [11]. We show that SAEs are a more suitable tool for interpretation and control as opposed to raw features.
SAEs for VLMs. Based on the success of SAEs to interpret features in LLMs, there have been several efforts in applying them to vision and vision-language models. Specifically, a popular direction has been to train SAEs on top of the CLIP model activations [35, 14, 36, 23] as well as other vision encoders (e.g. DINOv2 [31]). Additionally, there has also been interest in interpreting the denoising diffusion models using SAEs [37, 7, 19, 18], discovering common concepts across different vision encoders [39], and applying them on multimodal LLMs [47]. Differently, in this work, we focus on a rigorous evaluation framework for SAEs in VLM apart from utilizing SAEs on vision encoders to steer multimodal LLMs. Concurrent to our work, monosemanticity of multimodal models has been investigated [44], although with more focus on the difference between modalities. Our proposed MS score gives a more grounded quantitative measure of monosemanticity of individual neuron by taking into account their activations, while being aligned with human perception.
@@ 3.2 Monosemanticity Score
We formally describe our proposed MonoSemanticity score (MS) below, with an illustration given in Figure 2. This metric can be computed for each of the $\omega$ neurons extracted from the SAE. Given a diverse set of images =$\mathcal{I}=\{\mathbf{x}_{n}\in\mathbb{X}\}_{n=1}^{N}$, and a pretrained image encoder $E$, we first extract embeddings to obtain a pairwise similarity matrix $\mathbf{S}=[s_{nm}]_{n,m}\in[-1,1]^{N\times N}$, which captures semantic similarity between each pair of images. The similarity $s_{nm}$ of the pair $(\mathbf{x}_{n},\mathbf{x}_{m})$ is computed as the cosine similarity between the corresponding pair of embedding vectors:
@@
$\text{MS}^{k}:=\frac{1}{N(N-1)}\sum_{n=1}^{N}\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}r^{k}_{nm}s_{nm}.$ (9)
@@ 3.3 Steering MLLMs with Vision SAEs
Figure 3: Qualitative examples of highest activating images for different neurons from high (left) to low (right) MonoSemanticity scores (MS). As the score gets higher, images become more similar, illustrating the correlation with monosemanticity.
@@
$\forall j\in\{1,\dots,\omega\},\quad\hat{a}^{j}_{i}=\begin{cases}\alpha,&j=k\\ a^{j}_{i},&j\neq k\\ \end{cases}$ (10)
@@ 4.1 Experimental Settings
We train SAEs on ImageNet [9], to explain a fixed and pretrained CLIP-ViT L/14-336px [33] model, and a SigLIP model [45] in Appendix. Specifically, we pre-extract activation vectors from the explained model obtained on the images. The activation vectors are extracted from the classification (CLS) tokens in the residual stream after layers $l\in\{11,17,22,23\}$, or from the output of the (last) projection layer, except for steering experiments, in which we train SAEs from 2 random token embeddings per image extracted after layer $l=22$. In the following sections, we are interested in both BatchTopK [4] and Matryoshka BatchTopK SAEs [5] variants. If not stated otherwise, we set the groups of Matryoshka SAEs as =$\mathcal{M}=\{0.0625\omega,0.1875\omega,0.4375\omega,\omega\}$, which roughly corresponds to doubling the size of the number of neurons added with each level down. For the BatchTopK activation, we fix the maximum number of non-zero latent neurons to $K=20$. Both SAE types are compared across a wide range of expansion factors $\varepsilon\in\{1,2,4,8,16,64\}$. All SAEs are optimized for $10^{5}$ steps with minibatches of size $4096$ using Adam optimizer [21], with the learning rate initialized at $\frac{16}{125\sqrt{\omega}}$ following previous work [17]. To measure SAE performance, we use $R^{2}$ for reconstruction quality and the $L^{0}$ norm for the activation sparsity of $\phi(\mathbf{v})$. Throughout the paper, we quantify MS of neurons using DINOv2 ViT-B [31] as image encoder $E$, and present more analysis with different encoders in Appendix. Experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.2.1 Alignment of MS with human perception
Figure 4: Alignment rate (AR) (in %) of human answers with MS, with respect to the difference of the scores of the neurons associated to the highly activating images presented. Bars represent the AR within each interval, while dots indicate the cumulative AR up to and including that interval.
We first illustrate the correlation between MS score and the underlying monosemanticity of neurons, with examples in Figure 3 of the $16$ highest activating images for neurons with decreasing MS from left to right. We observe that the highest scoring neurons (with $\text{MS}=0.9$, on the far left) are firing for images representing the same object, i.e. close-up pictures of a koala (on the top) and a hockey (on the bottom). As the score decreases, the corresponding neurons fire for less similar or even completely different objects or scenes. To verify this observation in a more quantitative way, we conducted a large scale user study on the Mechanical Turk platform. This study resulted in a total of $1000$ questions across 71 unique users, with 3 answers per question aggregated through majority voting. Results of this study are presented in Figure 4, and we provide more details on the setup in Appendix. When asked to select the set of more monosemantic images from a pair of sets $(a,b)$, the users answered in accordance to the MS in $79.0\%$ of the cases, assuming $\delta=|\text{MS}(a)-\text{MS}(b)|\sim\mathcal{U}(0,0.9)$. This alignment rate monotonically raises from $55.7\%$ for $\delta\in(0.0,0.1)$ to $93.5\%$ for $\delta\in(0.8,0.9)$, highlighting that users especially agree with MS as the difference in monosemanticity between the two sets becomes more pronounced. This demonstrates that MS can be used as a reliable measure aligning well with human perception of similarity. Detailed results of the user study will be available for future benchmark of image similarity and monosemanticity.
@@ 4.2.2 Monosemanticity of SAEs
Table 1: Comparison of the highest Monosemanticity Scores (MS) of neurons from various layers of CLIP ViT-Large [33]. Sparse Autoencoders (SAEs) of two types decompose “No SAE” neurons into more monosemantic units. Higher SAE expansion factors yield higher MS of the resulting units.
@@
In Table 1, we report MS of the highest scoring neurons of two SAE types (BatchTopK [4] and Matryoshka BatchTopK [5]) trained at different layers with various expansion factors $\varepsilon$. We also include results for original neurons of the corresponding layer decomposed by SAEs (“No SAE”). We observe that SAEs’ neurons consistently have significantly higher MS for their best neuron when compared to original ones, implying that SAEs are better separating and disentangling concepts between their neurons. Interestingly, while the highest MS score is increasing with higher expansion factor $\varepsilon$, i.e. with increased width $\omega$ of the SAE layer, this holds true already for expansion factor $\varepsilon=1$, meaning that the disentanglement of concepts is also linked to the sparse dictionary learning and not only to the increased dimensionality. Finally, comparing SAE variants, we observe that while the Matryoshka reconstruction objective improves the concept separation at same expansion factor, it also achieves about 2 or 3 points lower $R^{2}$ for the same expansion factors (more details in Appendix).
(a) Impact of expension factor $\varepsilon$
@@
Figure 6: We steer the outputs of LLaVA by clamping the activation values of a chosen neuron, i.e. Neuron #39 = pencil neuron, in the CLIP SAE. We observe that while initially the poem follows the instruction (the prompt + white image) strongly, the outputs become more and more influenced by the concept that this neuron represents with increasing intervention weight $\alpha$, talking about attributes of pencil first, and then just the concept pencil. This shows that our interventions enable new capabilities for the unsupervised steering of these models.

### v3

@@ Abstract
Sparse Autoencoders (SAEs) have recently gained attention as a means to improve the interpretability and steerability of Large Language Models (LLMs), both of which are essential for AI safety. In this work, we extend the application of SAEs to Vision-Language Models (VLMs), such as CLIP, and introduce a comprehensive framework for evaluating monosemanticity at the neuron-level in visual representations. To ensure that our evaluation aligns with human perception, we propose a benchmark derived from a large-scale user study. Our experimental results reveal that SAEs trained on VLMs significantly enhance the monosemanticity of individual neurons, with sparsity and wide latents being the most influential factors. Further, we demonstrate that applying SAE interventions on CLIP’s vision encoder directly steers multimodal LLM outputs (e.g., LLaVA), without any modifications to the underlying language model. These findings emphasize the practicality and efficacy of SAEs as an unsupervised tool for enhancing both interpretability and control of VLMs. Code and benchmark data are available at https://github.com/ExplainableML/sae-for-vlm.
@@ 1 Introduction
In recent years, Vision-Language Models (VLMs) like CLIP [47] and SigLIP [61] have gained widespread adoption, owing to their capacity for simultaneous reasoning over visual and textual modalities. They have found a surge of applications in various modalities, such as in audio [14, 57] and medicine [62], transferring to new tasks with minimal supervision. Yet our current understanding of VLM internals remains limited, necessitating methods that can systematically probe their representations. Sparse AutoEncoders (SAEs) [37] are an effective approach to probing the internal representations of such models. They efficiently discover concepts (abstract features shared between data points) through their simple architecture learned as a post-hoc reconstruction task. Although analysis with SAEs is popular for Large Language Models (LLMs) [7, 48, 23], for VLMs it has been limited to interpretable classification [49, 35], or the discovery of concepts shared across models [53].
@@
Intuitively, SAEs reconstruct activations via a higher-dimensional space to disentangle distinct concepts from their overlapping representations in neural activations [7]. Neurons at different layers within deep neural networks are known to be naturally polysemantic [41], meaning that they can be strongly activated for multiple unrelated concepts such as cellphones and rulers. One common explanation for this behavior is the superposition hypothesis [3, 17], stating that concepts are encoded as linear combination of neurons. SAEs explicitly attempt to solve this issue by separating the entangled concepts into distinct representations. Despite their widespread use in research, the absence of a metric to evaluate SAEs at the neuron-level still hinders their practicality as an interpretation tool. Discovering neurons encoding human-interpretable concepts requires analyzing them individually, which is even more tedious as layers of both the model and SAE are wide.
@@
Figure 1: Sparse Autoencoder (SAE) in VLM (e.g. CLIP): Top activating images of a neuron in a pretrained VLM layer are polysemantic (left), and those of a neuron in a sparse latent of SAE trained to reconstruct the same layer are monosemantic (right), according to MonoSemanticity score (MS).
@@
In this work, we quantitatively evaluate SAEs for VLMs through monosemanticity, defined as the similarity between inputs that strongly activate a neuron. We propose the MonoSemanticity score (MS) for vision tasks, that measures the pairwise similarity of images weighted by the activations for a given neuron. Unlike natural language where individual words require surrounding context (such as complete sentences) to clearly identify their meanings, individual images can directly activate neurons without additional context. We define highly activating images as images that strongly fire a particular neuron. Greater similarity among these images suggests that the neuron is more narrowly focused on a single concept, reflecting higher neuron monosemanticity. Using our MS score, we observe and validate that neurons in SAE are significantly more monosemantic (see Figure 1, right) than the original neurons (see Figure  1, left). The neuron of the original VLM typically has a low MS score, it fires for a wide range of objects, from cellphones to rulers. On the other hand, neurons within the SAE are more focused on a single concept, e.g. parrots, obtaining a higher score. This holds even for SAE with the same width as the original layer, implying that the sparse reconstruction objective inherently improves the separability of concepts. We further conduct a large-scale study to quantitatively assess alignment of our proposed MS score with human interpretation of monosemanticity. The results confirm that the difference between scores of two neurons strongly correlates with humans perceiving the higher-scoring neuron as more focused on a single concept.
@@
Figure 2: Computation of our MonoSemanticity score (MS). (a) Embeddings and activations are extracted for a set of images (b) to compute the pairwise embedding similarities and pairwise neuron activations. (c) MS is the average of embedding similarities weighted by the neuron activations.
@@ 2 Related Work
Sparse Autoencoders. Recent studies have repurposed traditional dictionary learning to enhance LLM and VLM interpretability [6, 46]. Specifically, there has been success in interpreting and steering LLMs with features learned by SAEs [52, 16]. Several enhancements to SAE mechanisms have been introduced, including new activation functions such as Batch TopK [8] or JumpReLU [48], and insights from Matryoshka representation learning [34] for SAEs [39, 9]. We analyze and evaluate neuron-level monosemanticity of SAEs in VLMs and their downstream uses.
Vision-Language Models. Since Contrastive Language–Image Pretraining (CLIP) [47, 38, 61], many models have emerged that align images and text in a shared embedding space [56, 32] or generate text conditioned on image inputs [36, 12, 19]. They have achieved strong results on benchmarks [59] and found many use-cases  [24, 5, 26, 28]. As trust in these models has become a concern [60, 4, 27], understanding decision-making and ensuring safety, for example through steering, is increasingly important [29, 2]. Consequently, prior work has examined their internal representations [40, 21, 22, 45], uncovering interpretable neurons [15]. We demonstrate that SAEs enable more effective interpretation and control than directly operating on raw features.
SAEs for VLMs. Building on the success of SAEs in interpreting LLMs, researchers have tried applying them to vision and vision-language models, typically on CLIP [49, 20, 50, 35] and other vision encoders (e.g. DINOv2 [44]). There has also been interest in interpreting the denoising diffusion models [25] using SAEs [51, 11, 31, 30], discovering common concepts across different vision encoders [53], and applying them on multimodal LLMs [63]. Concurrently, monosemanticity of multimodal models has been investigated [58], although focusing on inter-modality differences. Our MS score provides a grounded, quantitative measure of the monosemanticity of individual neurons which is empirically aligned with human perception. Overall, we create a rigorous evaluation framework for SAEs in VLMs, as well as demonstrating their use for steering multimodal LLMs.
@@ 3.2 Monosemanticity Score
We formally describe our proposed MonoSemanticity score (MS) below, with an illustration given in Figure 2. This metric can be computed for each of the $\omega$ neurons extracted from the SAE. Given a diverse set of images $\mathcal{I}=\{\mathbf{x}_{n}\in\mathbb{X}\}_{n=1}^{N}$, and a pretrained image encoder $E$, we first extract embeddings to obtain a pairwise similarity matrix $\mathbf{S}=[s_{nm}]_{n,m}\in[-1,1]^{N\times N}$, which captures semantic similarity between each pair of images. The similarity $s_{nm}$ of the pair $(\mathbf{x}_{n},\mathbf{x}_{m})$ is computed as the cosine similarity between the corresponding pair of embedding vectors:
@@
$\text{MS}^{k}:=\frac{\sum_{1\leq n<m\leq N}r^{k}_{nm}\,s_{nm}}{\sum_{1\leq n<m\leq N}r^{k}_{nm}}$ (9)
@@ 3.3 Steering MLLMs with Vision SAEs
Figure 3: Top activating images of neurons with MonoSemanticity (MS) scores ranging from high (left) to low (right). Higher scores correlate with more similar images, reflecting monosemanticity.
@@
[…]
@@ 4.1 Experimental Settings
We apply SAEs to explain fixed and pretrained CLIP ViT-L/14-336px [47], SigLIP SoViT-400m/14-384px [61], AIMv2 L/14-224px [19], and WebSSL MAE-300m/14-224px [18]. The SAEs are trained on activation vectors pre-extracted from the model’s responses to ImageNet [13] images. For CLIP, activation vectors are extracted from the classification (CLS) tokens in the residual stream after layers $l\in\{11,17,22,23\}$, or from the output of the final projection layer. For steering experiments, however, the SAEs are trained on activation vectors corresponding to two random token embeddings per image, taken from layer $l=22$. For other encoders, we similarly use the CLS tokens from the final layers, or two random token embeddings if a CLS token is not available.
In the following sections, we are interested in both BatchTopK [8] and Matryoshka BatchTopK SAEs [9] variants. If not stated otherwise, we set the groups of Matryoshka SAEs as $\mathcal{M}=\{0.0625\omega,0.1875\omega,0.4375\omega,\omega\}$, which roughly corresponds to doubling the size of the number of neurons added with each level down. For the BatchTopK activation, we fix the maximum number of non-zero latent neurons to $K=20$. Both SAE types are compared across a wide range of expansion factors $\varepsilon\in\{1,2,4,8,16,64\}$. All SAEs are optimized for $10^{5}$ steps with minibatches of size $4096$ using Adam optimizer [33], with the learning rate initialized at $\frac{16}{125\sqrt{\omega}}$ following previous work [23]. To measure SAE performance, we use $R^{2}$ for reconstruction quality and the $L^{0}$ norm for the activation sparsity of $\phi(\mathbf{v})$. Throughout the paper, we quantify MS of neurons using DINOv2 ViT-B [44] as image encoder $E$, and present more analysis with different encoders in Appendix. Experiments are run on a single NVIDIA A100 GPU.
@@ 4.2.1 Alignment of MS with human perception
Figure 4: Alignment Rate (AR, %) of humans with MS score when judging which neuron in a pair is more monosemantic, grouped by MS difference between the neurons. Bars show AR per interval; dots show cumulative AR up to that interval.
We first illustrate the correlation between MS score and the underlying monosemanticity of neurons, with examples in Figure 3 of the $16$ highest activating images for neurons with decreasing MS from left to right. We observe that the highest scoring neurons (with $\text{MS}=0.9$, on the far left) are firing for images representing the same object, i.e. close-up pictures of a koala (on the top) and a hockey (on the bottom). As the score decreases, the corresponding neurons fire for less similar or even completely different objects or scenes. To verify this observation in a more quantitative way, we conducted a large scale user study on the Mechanical Turk platform. This study resulted in a total of $1000$ questions across 71 unique users, with 3 answers per question aggregated through majority voting. Results of this study are presented in Figure 4, and we provide more details on the setup in Appendix. When asked to select the set of more monosemantic images from a pair of sets $(a,b)$, the users answered in accordance to the MS in $82.8\%$ of the cases, assuming $\delta=|\text{MS}(a)-\text{MS}(b)|\sim\mathcal{U}(0,0.9)$. This alignment rate monotonically raises from $56.6\%$ for $\delta\in(0.0,0.1)$ to $100.0\%$ for $\delta\in(0.8,0.9)$, highlighting that users especially agree with MS as the difference in monosemanticity between the two sets becomes more pronounced. This demonstrates that MS can be used as a reliable measure aligning well with human perception of similarity. Detailed results of the user study are released for future benchmark of image similarity and monosemanticity.
@@ 4.2.2 Monosemanticity of SAEs
Table 1: Sparse Autoencoders (SAEs) decompose “No SAE” neurons into more monosemantic units as shown by MonoSemanticity (MS) score. Higher SAE expansion factors yield higher MS scores.
(a) Highest MS scores of neurons in various CLIP ViT-Large [47] layers.
@@
(b) Top MS scores of neurons from last layers of different vision encoders. Improvements in MS score from applying Matryoshka SAEs are consistent across all the models.
Vision Encoder No SAE Exp. factor
$\times$1 $\times$4
WebSSL [18] 0.01 0.79 0.92
CLIP [47] 0.01 0.82 0.89
SigLIP [61] 0.01 0.83 0.88
AIMv2 [19] 0.01 0.59 0.85
In Table 4(a), we report MS of the highest scoring neurons of two SAE types (BatchTopK [8] and Matryoshka BatchTopK [9]) trained at different layers with various expansion factors $\varepsilon$. We also include results for original neurons of the corresponding layer decomposed by SAEs (“No SAE”). We observe that SAEs’ neurons consistently have significantly higher MS for their best neuron when compared to original ones, implying that SAEs are better separating and disentangling concepts between their neurons. Interestingly, while the highest MS score is increasing with higher expansion factor $\varepsilon$, i.e. with increased width $\omega$ of the SAE layer, this holds true already for expansion factor $\varepsilon=1$, meaning that the disentanglement of concepts is also linked to the sparse dictionary learning and not only to the increased dimensionality. Finally, comparing SAE variants, we observe that while the Matryoshka reconstruction objective improves the concept separation at same expansion factor, it also achieves about 2 or 3 points lower $R^{2}$ for the same expansion factors (more details in Appendix).
(a) Impact of expansion factor $\varepsilon$
@@
Table 4(b) presents the highest observed MS scores among neurons in the last layers of various image encoders, both before (“No SAE”) and after attaching Matryoshka SAEs. We find that the SAE latent neurons outperform the original neurons in every case. As before, increasing the expansion factor $\varepsilon$ helps discover more monosemantic units. This suggests the universality of the SAE approach across vision representations derived from different training objectives.
@@
Table 2: Percentage of generations meeting evaluation criteria for concept insertion and suppression. SAE-derived steering directions yield higher success rates than Difference-in-Means (DiffMean) [2].
(a) Concept insertion
Ours DiffMean
Desired concept appeared 48.7 53.1
Base prompt followed 85.8 66.2
Both criteria satisfied 42.4 35.8
(b) Concept suppression
Ours DiffMean
Desired concept removed 64.4 64.0
Unrelated concept kept 81.4 38.7
Both criteria satisfied 52.5 33.3
Figure 6: Steered LLaVA outputs after clamping the activation values of a chosen neuron, i.e. Neuron #39 of SAE in CLIP corresponding to pencil. Initially, the poem follows the given instruction (the prompt and white image), but as the intervention weight $\alpha$ increases, it becomes increasingly influenced by the neuron’s concept, first mentioning the pencil’s attributes, then the pencil itself. This shows that our interventions enable new capabilities for the unsupervised steering of these models.